# Thousand Voices of Trauma: A Large-Scale Synthetic Dataset for Modeling Prolonged Exposure Therapy Conversations

### v3

@@ 4 Expert Evaluation of Synthetic PE Therapy Sessions
To ensure clinical relevance and gain insights into the utility and limitations of our synthetic dataset, we conducted an evaluation study involving clinical experts. Grounding synthetic data generation in real-world clinical perspectives is crucial for responsible development in AI mental health.
Seven therapists with diverse professional backgrounds (clinical practice, research, education) and extensive experience (6 to 30 years) across various settings (outpatient clinics, hospitals, VA/military, academia) evaluated two full synthetic PE therapy session transcripts. They assessed four key dimensions: content depth, perceived value, session appropriateness, and patient engagement, providing both quantitative ratings and qualitative feedback (See Figure 5).
The evaluation yielded valuable, multifaceted feedback. Experts consistently recognized the dataset’s strengths in capturing detailed patient narratives, with ratings ranging from “Somewhat detailed” (n=3) to “Very detailed” (n=4). Furthermore, the simulated patient’s engagement was rated positively, ranging from “Moderately” engaged (n=4) to “Extremely” engaged (n=3), indicating the model’s success in generating realistic and emotionally resonant patient responses crucial for PE simulations. Identified strengths included vivid trauma descriptions, significant emotional depth, and the portrayal of realistic novice therapist approaches – elements particularly useful for training applications. The feedback also highlighted areas reflecting the inherent challenges of synthetically generating complex therapeutic interactions.
Ratings for perceived value (ranging from “Not valuable” (n=2) to “Valuable” (n=1)) and appropriateness (from “Not appropriate” (n=3) to “Somewhat appropriate” (n=4)) varied, underscoring the difficulty in fully replicating nuanced clinical judgment and meeting all expert expectations synthetically. Similarly, the AI therapist’s perceived skill level (rated between “Novice” (n=2) and “Competent” (n=3)) suggests the current simulation captures less experienced therapist behavior more readily than expert-level interactions.
These critical assessments are invaluable, providing specific, actionable directions for future refinement. Key areas identified for improvement include enhancing conversational flow, reducing therapist interventions and repetitive prompts (e.g., “take a deep breath”), integrating more adaptive AI responses, and further increasing dialogue authenticity. This expert feedback highlights essential targets for advancing the sophistication of generative models in this sensitive domain.
Overall, the evaluation confirms that the LLM-generated transcripts effectively capture core elements of PE therapy sessions, particularly detailed patient narratives and recognizable therapeutic techniques. The diverse evaluator backgrounds provided robust, grounded feedback. While acknowledging the areas identified for continued refinement, this expert validation underscores the dataset’s immediate utility as a valuable resource for developing and testing AI models, studying specific therapeutic phenomena (like emotional arcs or novice therapist patterns), and providing a rich, privacy-preserving alternative to difficult-to-obtain real-world data. This rigorous evaluation process itself demonstrates a commitment to clinical fidelity and transparency in developing AI tools for mental health.
@@ 5.1 Rationale for a PE Therapy Benchmark
Developing a standardized benchmark is crucial for the responsible advancement and reliable comparison of AI models in trauma-focused care. PE therapy relies on carefully tracking and processing patient distress during imaginal exposure. AI tools intended to support PE delivery or therapist training must demonstrate fidelity to these core therapeutic dynamics. A consistent benchmark ensures potential AI applications align with clinical needs and therapeutic principles, facilitates reproducible research, and tracks progress in developing sophisticated AI for mental health support.
5.2 Benchmark Design
Our benchmark assesses emotional intensity across six conversational segments derived from each simulated therapy session, corresponding to different stages within PE therapy:
  1. 1.
    Orientation to Imaginal Exposure
  2. 2.
    Imaginal Exposure Duration
  3. 3.
    Monitoring SUDS Ratings
  4. 4.
    Reinforcing Comments
  5. 5.
    Eliciting Thoughts and Feelings (Processing Step 1)
  6. 6.
    Processing the Imaginal (Processing Step 2)
The expected emotional trajectory follows a recognizable pattern (Cowden Hindash et al., 2020; Bluett et al., 2014; Center for the Treatment and Study of Anxiety, 2023): initial anxiety during orientation, peaking distress during imaginal exposure, followed by gradual reduction through reinforcement and processing (see Figure 6). SUDS (Subjective Units of Distress Scale) are used to measure self-reported distress during exposure.
Although "Monitoring SUDS Ratings" is a distinct segment in our dataset, it occurs concurrently with the "Imaginal Exposure" phase in the PE protocol. Therefore, their emotional intensity ratings are expected to align closely, reflecting shared peak distress dynamics.
5.3 Evaluation Metrics
To quantify how well a model’s predicted emotional trajectory aligns with a baseline, we use three core metrics:
  • Sequence Similarity (Pearson Correlation): Measures the linear correlation between the model’s predicted emotional intensity sequence and the baseline sequence across the six phases. It assesses alignment in both magnitude and relative ranking of distress levels.
  • Pattern Accuracy (Dynamic Time Warping - DTW): Measures the similarity between the two sequences while allowing for non-linear warping in the time dimension. It captures how well the model follows the expected phase-wise progression, even with slight timing shifts. Lower DTW distances indicate better pattern accuracy.
  • Phase Consistency (Root Mean Squared Error - RMSE): Measures the average magnitude of error between the model’s predicted intensity score and the baseline score at each specific therapy phase. Lower RMSE indicates better phase-specific accuracy.
5.4 Baseline Establishment and Validation
@@
To establish a reference for comparison, we generated a Baseline Trajectory using zero-shot emotional intensity ratings from Claude Sonnet 3.5. Models assign scores on a 1 (calm) to 10 (extreme distress) scale for each of the six conversation segments per client profile. Claude Sonnet 3.5 was chosen for its strong zero-shot natural language understanding capabilities, suggesting suitability for interpreting emotional nuances without task-specific training. As shown in Figure 6, the resulting Baseline Trajectory aligns with established PE therapy patterns: peak distress occurs during the Imaginal Exposure and concurrent Monitoring SUDS Ratings phases, with gradual reductions during subsequent Reinforcing, Eliciting Thoughts/Feelings, and Processing phases. This correlates with the human evaluation and confirms the baseline’s suitability for evaluating future models.
5.5 Benchmark Evaluation Methodology
@@
We evaluated AI model performance by comparing their predicted emotional trajectories against the Claude Baseline Trajectory using the defined metrics (Pearson, DTW, RMSE), averaged across all 500 conversation profiles.
To compare models relative to the baseline, we compute performance ratios for each metric (M = model, C = Claude baseline):
@@
We selected a range of comparison models representing different sizes, architectures, and training methodologies (Mistral Large, Amazon Nova Pro, Llama3 70B/8B Instruct, Llama 3.1 70B/8B Instruct, Mistral 7B Instruct, Mistral Small) to test the benchmark’s ability to differentiate capabilities.
5.6 Results and Discussion
@@
The evaluation results, summarized in Table 1, demonstrate the benchmark’s ability to quantitatively differentiate model performance in assessing emotional trajectories.

### v4

@@ 4 Expert Evaluation of Synthetic PE Therapy Sessions
To ensure clinical relevance of our synthetic dataset, we conducted an evaluation study with seven therapists having diverse professional backgrounds (clinical practice, research, education) and extensive experience (6 to 30 years) across various settings (outpatient clinics, hospitals, VA/military, academia). They evaluated two full synthetic PE therapy transcripts, assessing content depth, perceived value, session appropriateness, and patient engagement (Figure 5). Experts recognized the dataset’s strengths in capturing detailed patient narratives, with ratings from “Somewhat detailed” (n=3) to “Very detailed” (n=4). The simulated patient’s engagement was rated positively from “Moderately” engaged (n=4) to “Extremely” engaged (n=3), indicating the model’s success in generating realistic, emotionally resonant responses crucial for PE simulations. Strengths included vivid trauma descriptions, emotional depth, and realistic novice therapist approaches—elements particularly valuable for training applications. However, ratings for perceived value (from “Not valuable” (n=2) to “Valuable” (n=1)) and appropriateness (from “Not appropriate” (n=3) to “Somewhat appropriate” (n=4)) varied, highlighting challenges in synthetically replicating nuanced clinical judgment. The AI therapist’s skill level (rated between “Novice” (n=2) and “Competent” (n=3)) suggests the simulation better captures less experienced therapist behavior than expert-level interactions. Key improvement areas include enhancing conversational flow, reducing repetitive interventions (e.g., “take a deep breath”), developing more adaptive AI responses, and increasing dialogue authenticity. Despite these limitations, the evaluation confirms the LLM-generated transcripts effectively capture core elements of PE therapy—particularly detailed patient narratives and recognizable therapeutic techniques. This expert validation underscores the dataset’s utility as a valuable resource for developing AI models, studying specific therapeutic phenomena, and providing a privacy-preserving alternative to difficult-to-obtain real-world data. A companion study compares these synthetic sessions with real PE transcripts, showing strong structural and emotional alignment [34].
To evaluate whether AI models can meaningfully interpret the emotional progression in synthetic PE sessions, we design a benchmark focused on emotional trajectory fidelity.
@@ 5 Benchmark Setup and Evaluation
Rationale for a PE Therapy Benchmark: Developing a standardized benchmark is crucial for the responsible advancement and reliable comparison of AI models in trauma-focused care. PE therapy relies on carefully tracking and processing patient distress during imaginal exposure. AI tools intended to support PE delivery or therapist training must demonstrate fidelity to these core therapeutic dynamics. A consistent benchmark ensures potential AI applications align with clinical needs and therapeutic principles, facilitates reproducible research, and tracks progress in developing sophisticated AI for mental health support.
Benchmark Design: Our benchmark assesses emotional intensity across six conversational segments derived from each simulated therapy session, corresponding to different stages within PE therapy: (1) Orientation to Imaginal Exposure, (2) Imaginal Exposure Duration, (3) Monitoring SUDS Ratings, (4) Reinforcing Comments, (5) Eliciting Thoughts and Feelings (Processing Step 1), and, (6) Processing the Imaginal (Processing Step 2). The expected emotional trajectory follows a recognizable pattern [35, 36, 37]: initial anxiety during orientation, peaking distress during imaginal exposure, followed by gradual reduction through reinforcement and processing (see Figure 6). SUDS (Subjective Units of Distress Scale) are used to measure self-reported distress during exposure. Although “Monitoring SUDS Ratings” is a distinct segment in our dataset, it occurs concurrently with the “Imaginal Exposure” phase in the PE protocol. Therefore, their emotional intensity ratings are expected to align closely, reflecting shared peak distress dynamics.
Evaluation Metrics: We evaluate alignment between a model’s predicted emotional trajectory and the baseline using three metrics: Sequence Similarity (Pearson Correlation): captures linear correlation across the six phases, reflecting agreement in both magnitude and distress ranking; Pattern Accuracy (DTW): applies Dynamic Time Warping to account for temporal flexibility, with lower values indicating better phase-wise adherence; and Phase Consistency (RMSE): computes the average per-phase error, where lower values reflect higher accuracy.
@@
Baseline Establishment and Validation: We establish a Baseline Trajectory using zero-shot emotional intensity ratings from Claude Sonnet 3.5, selected for its strong general language understanding and ability to interpret emotional nuance without task-specific training. The model rates each of six conversation segments per client profile on a 1 (calm) to 10 (extreme distress) scale. As shown in Figure 6, the resulting trajectory mirrors expected PE therapy patterns: peak distress during Imaginal Exposure and Monitoring SUDS Ratings, followed by gradual declines through Reinforcing, Eliciting Thoughts/Feelings, and Processing. This alignment with human evaluations validates its use as a reference for future models.
@@
Benchmark Evaluation Methodology: We evaluated AI model performance by comparing their predicted emotional trajectories against the Claude Baseline Trajectory using the defined metrics (Pearson, DTW, RMSE), averaged across all 500 conversation profiles. To compare models relative to the baseline, we compute performance ratios for each metric (M = model, C = Claude baseline):
@@
We selected a range of comparison models representing different sizes, architectures, and training methodologies (Mistral Large, Amazon Nova Pro, Llama3 70B/8B Instruct, Llama 3.1 70B/8B Instruct, Mistral 7B Instruct, Mistral Small) to test the benchmark’s ability to differentiate capabilities. All models were accessed via Bedrock API, and we use Sonnet 3.5 both for dataset generation and as a zero-shot baseline for evaluation.
@@
Results and Discussion: The evaluation results, summarized in Table 1, demonstrate the benchmark’s ability to quantitatively differentiate model performance in assessing emotional trajectories.